OpenStudy (amenah8):

write the equations in slope intercept and standard form: perpendicular to 3x+4y=13 and through (2,7) parallel to y= -5 and through (2,7) vertical and contains (-4,6) x-intercept= 2 and y-intercept= -4

OpenStudy (triciaal):

slope intercept form y = mx + b for any point on the line (x,y) m = slope and b= y-intercept (when x = 0) standard form Ax + By = C A,B and C are real numbers parallel lines have the same slope for perpendicular lines the product of the slopes is -1 now go through each problem with this in mind

OpenStudy (amenah8):

ok, thanks. but for the parallel problem, there is no slope. so i would have to solve for slope first right?

OpenStudy (amenah8):

and are you saying that for all perpendicular lines, the slope in -1?

OpenStudy (triciaal):

no! the product of the slopes =-1

OpenStudy (triciaal):

slope of line AB multiplied by slope of line PQ = -1 for lines AB and PQ

OpenStudy (amenah8):

i still don't understand how i would solve the first problem with the information i am given. i know that the 3x+4y=13 is in standard form. but what would i do to get the standard form of the line it is perpendicular to??

OpenStudy (triciaal):

m1*m2 = -1 where m1 is one slope and m2 is the slope of the perpendicular line

OpenStudy (amenah8):

so how would i find the slope?

OpenStudy (triciaal):

first rewrite to y = mx + b form m = slope let this slope = m1 the perpendicular let it be = m2 use m1*m2 = -1 to find m2 you have a point on the perpendicular line find b then write the general eqn

OpenStudy (amenah8):

i think i understand now. thank you so much!!

OpenStudy (triciaal):

great welcome

OpenStudy (amenah8):

how would i solve for b??

OpenStudy (triciaal):

use the point on the line (x.y) (2,7)

OpenStudy (triciaal):

y = mx + b

OpenStudy (amenah8):

thanks!

OpenStudy (triciaal):

let me see when you have it in standard form

OpenStudy (amenah8):

ok

OpenStudy (amenah8):

b would be 7, right?

OpenStudy (amenah8):

x+4y=28 is standard form

OpenStudy (triciaal):

that's not what i got

OpenStudy (amenah8):

b is 7, right??

OpenStudy (triciaal):